eah3:

A retired woman has $50,000 to invest but needs to make $6,000 a year from the interest to meet certain living expenses. One bond investment pays 15% annual interest. The rest of it she wants to put in a CD that pays 7%. Set up and solve the equation for how much the woman should invest in each option to sustain exactly a $6,000 annual return

Hero:

Let x represent the amount of money in one bond and y represent the amount of money in the certificate of deposit Then we can set up the following system: x + y = 50000 0.15x + 0.07y = 6000 After some manipulation we have: 15x + 7y = 600000 7x + 7y = 350000 After subtraction we have: 8x = 250000 Solving for x and y yields: x = 31250 y = 18750 So $31,250 should be invested in the bond and $18,750 should be invested in the certificate of deposit. Let's check this: 31250 + 18750 = 50000 .15(31250) + .07(18750) = 6000 4687.50 + 1312.50 = 6000

Hero:

@psirockin2 do the amounts you come up with sum to $50,000